Mathematics 244 – Linear Algebra Definitions Quiz 1 – January 26, 2007

Your Name:

A) (5) What is the definition of the span of a set of vectors in R^{n}?

B) (5) The matrix A below can be into reduced row echelon form (rref) using exactly two row operations. What are they and what is the rref?

( \begin{matrix}2 & −1 & 0 & 4 & 0 & −3 \\ 0 & 0 & 0 & 0 & 1 & 5 \\ 0 & 0 & 1 & −2 & 0 & 7 \end{matrix} )